in numericals related to relative lowering of vapour pressure,sometimes ?p/p =x(mole fraction of solute) is used & sometimes
?p/p =n2/n1 (n2=moles of solute,n1=moles of solvent)is used.I know second formula is used in case of dilute solution but still I am not able to distinguish their use.please help me

Grade:

1 Answers

Mujahid Ahmed
11 years ago

You are mixing up the formulas

(P0-p)/p = NA/NB

(P0-p)/p0 = Xa

Focus on what the denominator is.

 

Approve if i helped :)